# Victoria, BC, Canada

### v1

@@ 4 A Landscape of Mixed and Multi Methods Design
Research study design that involves choosing suitable procedures and deciding on their timing is a challenging process easterbrook_selecting_2008 , but research design is particularly challenging in the case of an MMMR study. The research problem or questions being asked and how they emerge over the course of a study guides the research design and the timing of procedures used. Below we discuss the main characteristics of a research design, followed by some common patterns of MMMR designs.
4.1 Core Characteristics of Mixed and Multiple Research Designs
@@
  • Planned or Emergent Design: A mixed or multiple research design may be planned up front as guided by the research problem or question, or it may organically evolve over time because a first approach may not work as well as anticipated (for example, due to low participation), or it may lead to insights that beg for confirmation, expansion, explanation, or triangulation.
@@
  • Inductive/Deductive Dominance: An MMMR design typically involves both inductive (from data to theory) and deductive research (from theory to data), but one approach usually dominates or takes priority over the other driven by the philosophical world view of the researcher or by the nature of the research problem or question. Researchers often conflate inductive and deductive research with quaL and quaN methods, respectively, but Young et al. explain how inductive and deductive research may fall along a “spectrum” of quaL and quaN data young_spectrum_2020 .
@@
  • Mixed or Multi Method Design: When the research mixes quaL and quaN data (during collection or analysis), the design is referred to as mixed rather than a multi methods design, which uses both quaL or quaN procedures.
  • Data Integration: How the data is integrated during collection or analysis shapes the research design and the research insights ladner_mixed_2019 (as mentioned above). For example, in the case of parallel MMMR, quaN and quaL data may be merged to lead to the findings or insights, or in the case of a sequential study, one type of data may build on top of or add to another type (see Figure 1).
@@ 4.2 Research Designs
The characteristics mentioned above and how they are composed lead to a constellation of different research designs. We use the nomenclature for research designs as proposed by Creswell and Plano Clark creswell_best_2011 as we describe some popular design variations in the following.
@@
  • Exploratory Sequential: This sequential design often prioritizes a quaL method (normally the dominant method) followed by a quaN procedure that builds on insights from the first procedure. The design may be planned up front or it may emerge as insights are uncovered from the first procedure.
@@
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first procedure. The design may be planned up front or the need for the second procedure may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation.
@@
  • Convergent Parallel: The convergent parallel design involves both a quaL procedure and a quaN procedure to bring complementary insights on a research question. The goal may be to uncover both convergent and divergent insights from both procedures that are planned in parallel.
@@
  • Embedded: An embedded or nested design is when one procedure is dominant (quaL or quaN ) but another procedure appears as a “strand” within the first procedure. For example, as part of a quaN survey, a researcher may choose to interview select participants while carrying out the survey to gain more quaL insights on their answers.
  • Multi-method: A multi-method design may be planned or emergent and involves either two or more quaN procedures or two or more quaL procedures, but not a mix of both quaL and quaN. The main goal is typically to expand on the insights across procedures, and they may be carried out sequentially or in parallel.
5 Researcher Personas and MMMR Design Scenarios
To illustrate the principles, characteristics and MMMR designs we presented so far in this paper, we showcase four fictitious personas as guiding examples. These four fictional software engineering researchers are at different stages in their research careers, pursuing quite different research projects, each with a different reason why a mixed or multi methods design is appropriate. We describe these below and mention the decisions and challenges they faced in conducting their research. For the four scenarios, we provide illustrations of the research designs (see Figures 3, 4, 5 and 6). Later we summarize the four scenarios in Table 1.
5.1 Ali’s Study: The Impact of Security Experts Assigned to Development Teams
@@
Figure 3: Ali follows an exploratory sequential design, studying the impact of security experts on development practices. The unit of analysis is practices as informed by developers across both procedures.
Ali (he, him), is a postdoc studying the introduction of a new process whereby security experts are assigned to development teams to improve the compliance of security practices by developers at a large company. Ali’s research question is to ask if this change impacts developer motivations and capabilities for writing more secure code.
@@
As this is a new process and there is little research to build on, Ali first conducts an inductive exploratory open-ended interview with a small number of developers at a company to find out how this change has impacted developer motivations and behaviours for writing more secure code. As Ali conducted these interviews, he took care to protect the identity of the team members he interviewed, with both the security and ethical concerns in mind. The findings from these interviews help identify a number of concepts that help describe changes in motivation.
Ali wishes to know if the interview findings from this team hold (i.e., generalize) across a broader set of developers. He expands the preliminary exploratory approach and uses the insights from the initial interviews to develop a follow-up closed question validating survey with a larger population of developers beyond the company to confirm if positive changes in motivation occur when security experts are added to software teams. Again, the identity of the developers was protected and care was taken to ensure the presentation of the findings neither compromises the developers’ roles nor the confidentiality of the projects they work on.
Both methods were conducted in a rigorous way and led to novel insights about the impact of security experts on developer teams and how those hold across a larger population of developers. Ali planned to have both studies at the start of the research and sees the qualitative insights as the dominant contribution in this exploratory sequential research design. The tradeoff Ali made in this design is to focus some of his research effort on the follow-up survey to generalize the findings over achieving potentially deeper insights from longer or more interviews in the first phase.
5.2 Sam’s Study: The Impact of a New Code Review Tool on Code Quality
@@
Figure 4: Sam follows an explanatory sequential design to study a new code review tool. The unit of analysis is telemetry process metrics in (1), while in (2) it is perception of the process informed by developers.
@@
Sam (they, them) is an early stage Ph.D. student collaborating with a small company that is trialing a new tool to automate part of the code review process with hypothesized improvements on time-to-merge of code changes and on the quality of the reviewed code. They start their research with a deductive quantitative data mining study to investigate their hypothesis that this new tool (before and after use) improves time-to-merge and reduces the number of bugs captured in their industry partner’s code repository. As they mine the data, Sam ensures that their findings do not reveal whose coding or code review activities may have resulted in the future introduction of bugs with other team members or leads.
@@
From the mining study phase, Sam finds that the tool results in the hypothesized faster time-to-merge but not a lower number of bugs after merge. This insight was not expected, thus to understand why, they then conduct open ended interviews with a select number of developers to explore how the tool was used and to investigate inductively why it did not have the hypothesized impact on bug count. Again, care is taken to protect the identity of the developers interviewed, but also to ensure the interviews did not make them feel stressed given the presence of some bugs they may have introduced. From the interviews as well as the mined data, novel insights about the tool emerged.
@@
The use of mixed methods was not intended up front, but was an emergent or organic research design decision initiated by the surprising findings. These integrated insights can be used to shape code review tool design accordingly. The resulting design is referred to as an explanatory sequential design as mentioned above (see Figure  4). A tradeoff was made to halt further mining of the telemetry data and instead to conduct interviews with developers to explain the surprising results, but also to shed light on more actionable insights for how the code review tool could be improved.
5.3 Vicki’s Study: The Adoption of Generative AI at a Software Company
Figure 5: Vicki follows an emergent convergent parallel design to study the adoption of Generative AI. The unit of analysis is code changes by directly observing code.
Vicki (she, her) is an experienced industry researcher working at a software company, and her research aim is to understand the grassroots adoption of generative AI technologies across her company. As this technology is so new, and not much is known about how it can and will be used to enhance developer productivity, she decides to inductively consider the AI adoption from two angles in parallel. First, Vicki uses an ethnographic approach of observations to glean how the generative AI tool is used by developers at her organization. She observes its adoption on her own team of developers, while taking care of ethical concerns by anonymizing insights on individual productivity as the use of a new tool like AI may stress developers if they worry about losing their jobs.
In parallel to complement these observations, she carefully mines logging data from the AI tool used by multiple teams and in projects at her organization to understand more about the scale of adoption and how the tool is influencing developer productivity in terms of number of commits and speed of feature development. She recognizes that mining this data comes with the potential risk of introducing privacy concerns as she works at the company, but she mines in a way to obfuscate developers’ identities and does this to confirm her observations (which could be biased by her own impressions).
The use of a convergent parallel design that involves conducting both procedures in parallel provides complementary and novel insights on the adoption of this new potentially disruptive technology at her company.
@@ 5.3 Vicki’s Study: The Adoption of Generative AI at a Software Company
5.4 Zara’s Study: Understanding the Effectiveness of an Automated tool for Generating Tests
@@
Figure 6: Zara follows multi-method design to study a test generation tool. The unit of analysis is the developers in (1), while in (2) it is the test cases generated by the tool.
@@
Zara (she, her) is a late stage Ph.D. student who aims to investigate how effective an automated tool she designed is for generating tests. Before asking practitioners, she initially conducts a lab experiment with senior student developers, where half use the tool and half do not. She takes care during recruitment (via a university professor) to be cautious about how she recruits the students.
She starts deductively with the hypothesis that the tool will help with faster test generation. She assigns the same testing tasks to both groups. Through self reports from the developers during the experiment, Zara finds that the student developers report that through the use of her tool they are faster at generating test cases, supporting her hypothesis and also revealing that they believe the tool makes them faster.
This positive finding encourages her to expand her research question to investigate if the tests generated by her tool are as or more effective at finding bugs. She follows up by mining test and bug data generated during the lab experiment to investigate if the tests generated are effective at defect localization. Although she is aware of the identity of the students from the experiment, she takes care to obfuscate their data when she mines it.
Note she expands not only the research question she is asking, but she also changes the unit of observation in her study (from observing developers to analyzing the test cases generated). In this case, her research design is a multiple methods design rather than a mixed design as she does not mix qualitative and quantitative data while both of her studies are deductive by nature. Following her study, she has gained novel insights on how students may benefit from the test generation tool. An alternative design to this multiple methods design could have been to only mine data from the students’ activities. However, she used two methods as she was also interested to gain insights on their perception of the tool’s effectiveness.
Type of Design Design flow Rationale(s) Persona
@@ 3. Vicki
Generative AI
Multi-Method Design
@@
QuaN: Deductive Data mining
@@ 4. Zara
Table 1: The first three scenarios are mixed method designs, while the fourth is a multiple methods design. Note that there may be several reasons for conducting a mixed or multiple method design, and this may not be apparent up front (e.g., in the case of Sam).
@@ 5.5 Learning from Other Examples
So far, we have shared that using MMMR enables SE researchers to study their topics from a more holistic perspective, expanding and complimenting methods to achieve more and novel insights. Furthermore, we provide personas and scenarios to demonstrate the broad and varied landscape of possible MMMR designs in SE research. The ACM SigSoft Empirical Standards for Multiple and Mixed Methods also provides a list of example MMMR papers sigsoft , but care has to be taken in using these and other examples of MMMR designs as “exemplars” for future work. With this in mind, we next present “antipatterns” to watch out for when designing or critiquing an MMMR design. An earlier version of these antipatterns also appears in the same Standards, contributed by the first author of this paper.
6 Antipatterns of Mixed and Multi Method Research Designs
The principles and research design patterns we described above can be considered as useful guidelines to follow when designing a study. And through the scenarios above, we illustrated how the principles of MMMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers and reviewers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented (rather than inherent to the conduct of the research). We introduce the presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be spotted at design time (or during review of a paper). We recommend researchers think about these antipatterns early in their research process.

### v2

@@ 4 A Landscape of Mixed Methods Design
Research study design that involves choosing suitable procedures and deciding on their timing is a challenging process easterbrook_selecting_2008 , but research design is particularly challenging in the case of an MMR study. The research problem or questions being asked and how they emerge over the course of a study guides the research design and the timing of procedures used. Below we discuss the main characteristics of a research design, followed by some common patterns of MMR designs.
4.1 Core Characteristics of Mixed Research Designs
@@
  • Planned or Emergent Design: A mixed research design may be planned up front as guided by the research problem or question, or it may organically evolve over time because a first approach may not work as well as anticipated (for example, due to low participation), or it may lead to insights that beg for confirmation, expansion, explanation, or triangulation.
@@
  • Inductive/Deductive Dominance: An MMR design typically involves both inductive (from data to theory) and deductive research (from theory to data), but one approach usually dominates or takes priority over the other driven by the philosophical world view of the researcher or by the nature of the research problem or question. Researchers often conflate inductive and deductive research with quaL and quaN methods, respectively, but Young et al. explain how inductive and deductive research may fall along a “spectrum” of quaL and quaN data young_spectrum_2020 .
@@
  • Data Integration: How the data is integrated during collection or analysis shapes the research design and the research insights ladner_mixed_2019 (as mentioned above). For example, in the case of parallel MMR, quaN and quaL data may be merged to lead to the findings or insights, or in the case of a sequential study, one type of data may build on top of or add to another type (see Figure 1).
@@ 4.2 Research Designs
The characteristics mentioned above and how they are composed lead to a constellation of different research designs. We use the nomenclature for research designs as proposed by Creswell and Plano Clark creswell_best_2011 while describing some popular design variations in the following. As we describe each design variation, we provide an example of a recent paper published in a relevant empirical software engineering journal that uses that design. [^7] We found looking through the last twelve issues of the Empirical Software Engineering journal and the Information and Software Technology journal, that many SE papers use mixed methods, but they often do not say they are mixed methods research, and they also usually do not articulate or justify the specific design they use. Here, we provide one example paper that maps to each research design. If the paper was not that clear about the design, we contacted the authors to confirm our mapping.
@@
  • Exploratory Sequential: This sequential design often prioritizes a quaL method (normally the dominant method) followed by a quaN procedure that builds on insights from the first procedure. The design may be planned up front or it may emerge as insights are uncovered from the first procedure. For example, Alami et al. Alami2024 explore how psychological safety impacts agile team behavior through a mixed method approach consisting of an initial exploratory phase with 20 interviews, followed by a quaN survey study involving 423 participants.
@@
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first procedure. The design may be planned up front or the need for the second procedure may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation. For example, Jiang et al. Jiang2024 mined open source deep learning projects to analyze the distribution of computer vision defects, followed by interviews conducted with leaders of a reengineering team to shed light on the challenges and practices that may have led to the defects.
@@
  • Convergent Parallel: The convergent parallel design involves both a quaL procedure and a quaN procedure to bring complementary insights on a research question. The goal may be to uncover both convergent and divergent insights from both procedures that are planned in parallel. For example, Hidellaarachchi et al. Hidellaarachchi2024 carry out personality test-based surveys and in-depth interviews with a selected set of participants to investigate the impact of personality on requirements engineering activities. The survey and interviews were conducted in parallel and insights obtained from both approaches were used to answer the research questions.
@@
  • Embedded: An embedded or nested design is when one procedure is dominant (quaL or quaN) but another procedure appears as a “strand” within the first procedure. For example, as part of a quaN survey, a researcher may choose to interview select participants while carrying out the survey to gain more quaL insights on their answers. A representative paper that describes using this research design can be found in the work by Madugalla et al. Madugalla2024 where the mixed (quaL+quaN) survey procedure is dominant, followed by nested (quaL) interviews. Their study aimed to identify challenges and possible benefits of conducting SE research with human participants during the pandemic. It nested the findings of mixed survey data with in-depth follow up interviews with those researchers who had answered the survey and volunteered for the follow up interviews.
5 MMR Design Scenarios
To illustrate the principles, characteristics and MMR designs we presented so far in this paper, we showcase four scenarios as guiding examples. There could be many more study design examples, but we focus on four that map to each of four MMR designs (exploratory sequential, explanatory sequential, convergent parallel, and embedded) for the sake of brevity and clarity. For each scenario, we suggest a fictional SE researcher who is pursuing a mixed method design, and use these characters to show how having empathy and understanding for our study participants can also lead to more thoughtful and rigorous research.
We describe the four scenarios below and mention the decisions and challenges each fictional researcher could have faced in designing and conducting their research. For the four scenarios, we provide illustrations of the research designs (see Figures 3, 4, 5 and 6). We also summarize the four scenarios in Table 1.
5.1 Ali’s Study: Investigating the impact of a new security tool on developer productivity and developer experience at a large software company.
@@ 5 MMR Design Scenarios
Figure 3: Ali follows an exploratory sequential design, studying the impact of a security tool that enforces secure development practices. The unit of analysis are the practices informed by the developers using the tool across a company.
Ali (he, him), is a postdoc working closely with a partner company studying the introduction of a new security tool that is used by development teams. The use of the tool enforces the compliance of security practices by developers at this large company. Introducing a new tool like this that enforces changes in development work has complex impacts not just on the nature of work, but also on the developers’ lived experience. Ali’s research question is to investigate how this tool impacts perceived developer productivity as well as their experience while developing. Ali has prior industry experience and recognizes the complexity of introducing this new tool and decides to use mixed methods to bring and integrate insights on the technical and social aspects of the new tool.
@@
As Ali finds little related research to build on, he first conducts an inductive exploratory open-ended interview with a small number of developers from a team at his partner (case) company to find out how this change has impacted their developer practices for writing more secure code, and also how it has impacted their perceived productivity and experience. As Ali conducts these interviews, he takes care to protect the identity of the team members he interviewed, with both the security of the organization and ethical concerns with respect to participant job security in mind. For example, developers may be frustrated by such a tool but may not want to openly share this at their company. The findings from these interviews helps identify a number of concepts that reveal changes in developer productivity and experience, in addition to changes in the perceived quality and security of the code.
Ali finds that, contrary to his expectations from his prior industrial experience with mandated security tools, the tool does not negatively impact perceived developer productivity nor does it negatively impact their experience. He wishes to know if the interview findings from the members on this team hold (generalize) across a broader set of developers at his partner company. He also learned that some developers changed the tools they were using, a finding he did not expect. He expands the preliminary exploratory approach and uses the insights from the initial interviews to develop a follow-up closed question survey with a larger population of developers at the same company. In addition to asking questions about their perceived productivity, and experience, he also asks additional questions about the tools they use, an expansion over what he intended originally. In the survey, the identity of the developers and their projects is again protected.
Both methods were conducted in a rigorous way and led to novel insights about the impact of a security compliance tool on perceived developer productivity and their experience. Through the survey he also learned answered some questions prompted by the interviews across a larger population of developers. Ali wanted to conduct both procedures (sequentially) at the start of the research and considers the qualitative insights as the dominant contribution in this exploratory sequential research design.
The trade-off Ali makes in this design is to focus some of his research effort on the follow-up survey to triangulate, to generalize, the findings over achieving potentially deeper insights from longer or more interviews in the first phase. But the benefits of using both approaches is that he gains insightful results and finds they hold across more teams and in different team contexts. See Figure 3 for an overview of Ali’s research approach.
5.2 Sam’s Study: Studying The Impact of a Generative AI Tool that Automates Coding Reviewing
@@
Figure 4: Sam follows an explanatory sequential design to study the adoption of a generative AI reviewing tool on code quality. The unit of analysis is telemetry process metrics in (1), while in (2) it is perception of the process informed by developers.
Sam (they, them) is an early stage Ph.D. student collaborating with a large company that is trialing a new Generative AI tool to automate part of the code review process with hypothesized improvements on time-to-merge of code changes and on the quality of the reviewed code.
@@
Sam starts their research with a deductive quantitative data mining procedure to investigate their hypothesis that this new tool (before and after use) improves time-to-merge and reduces the number of bugs captured in their industry partner’s code repository. As they mine the data, Sam ensures that their findings do not reveal whose coding or code review activities may have resulted in the future introduction of bugs with other team members or leads.
From the mining of data, Sam confirms the tools lead to the hypothesized faster time-to-merge but not a lower number of bugs after merge. Sam wishes to understand this unexpected result, and so to gather the views of the developers who used this tool, Sam conducts open ended interviews with a random selection of developers to probe inductively why it did not have the hypothesized impact on bug count.
@@
Care is taken to protect the identity of the developers interviewed, but also to ensure the interviews did not make participants feel stressed given the presence of some bugs they may have introduced. From the interviews, as well as the mined data, novel insights about the tool and how it was used, emerged. Notably, from the interviews, Sam learns that the developers are able to work more efficiently but they are perhaps a little more complacent using the AI reviewing tool.
The use of mixed methods was not intended up front, but was an emergent or organic research design decision so that an explanation could be found through the interviews. The integrated insights can be used to shape code review tool design accordingly, by adding features to prompt developers to avoid complacency.
@@
The resulting research design Sam followed is an explanatory sequential design (see Figure 4). A trade-off was made to halt further mining of the telemetry data indicating speed of merge and number of bugs, to instead conduct interviews with developers to explain the surprising results, but also to shed light on more actionable insights for how the code review tool could be improved.
5.3 Vicki’s Study: Understanding the Adoption of a New Onboarding Process in a Large Scale Open Source Project
Figure 5: Vicki follows a planned convergent parallel design to study the adoption of a new onboarding process in an open source project.
Vicki (she, her) is an experienced researcher that has been studying collaboration practices in open source projects for some time. Her most recent research aim is to understand the adoption of a new onboarding process to address the retention of new contributors to an open source project that she has been involved with (as a contributor and researcher) for some time.
The onboarding process relies on a form of mentoring that has not been studied before. This process suggests that seasoned and new developers pair on work items, asynchronously communicating over Slack).
She decides to inductively consider the onboarding process from two angles in parallel. Vicki uses observations of their communication over the project’s asynchronous communication channel to discern how the onboarding process is used by both existing and new contributors in the project. See Figure 5 for an overview of her research design. Note, because of ethical concerns, Vicki anonymizes the identity of the participants she observes.
In parallel to, and to complement her observations, she mines logged data about code commits from the new contributions that are made by the new contributors. She studies the quality of their first commits but also considers the timing and quality of their subsequent commits.
She recognizes that mining the commit data comes with the potential risk of introducing privacy concerns, so she obfuscates developers’ identities. She finds that mentoring during the onboarding process leads to higher quality commits and an increase in commits for those who were mentored.
The use of a convergent parallel design that involves conducting both qualitative and quantitative procedures in parallel provides complementary and novel insights on the adoption of this new onboarding process in this open source community. She finds that the onboarding mentorship process of pairing and communicating over Slack leads to an increase in future contributions and higher quality commits.
@@
5.4 Zara’s Study: Evaluating the Effectiveness of an Automated tool for Generating Tests
@@
Figure 6: Zara follows a mixed methods embedded design to study a test generation tool. The unit of analysis is the test cases generated by the tool, in addition to the experience of the participant using (or not using) the automated tool.
@@
Zara (she, her) is a late stage Ph.D. student who aims to investigate how effective an automated tool she designed is for generating tests, in comparison to not using the tool. She initially conducts a lab experiment with senior student developers, where half use the tool and half do not. She takes care during recruitment that the students’ professor is not involved and thus may coerce students to participate (an ethical concern).
She starts, deductively, with the hypothesis that the tool will lead to faster test generation. She assigns the same testing tasks to both groups. Zara finds that the student developers are faster at writing tests using the automated tool. For a random selection of the participants, thinking inductively, she asks the students to qualitatively describe their confidence about the tests they created, and how the tool impacts their knowledge about testing in general, through intermittent short surveys they fill out between tasks. She finds from the two methods, that the students using the automated tool are not only faster, but they also report more confidence in their tests and learned more about testing.
An alternative design to the mixed methods design could have been to mine data from the students’ activities as a single method. However, she used two mixed methods as she was also interested to gain insights on how the tool impacted their confidence and learning about testing. Another advantage of using the embedded surveys, is that the surveys also helped explain, and lead credibility to, the quantitative results from the mined data.
The embedded design she used is primarily a quantitative method, examining the speed of tests created, but has an embedded qualitative component, to find out additional insights about their perceived confidence and learning (see Figure 6 for an overview of this MMR design).
Type of Design Design flow Rationale(s) Scenario
@@ 3. Vicki
Onboarding Mentorship
Embedded
@@
quaL: Surveys
@@ 4. Zara
Table 1: Four fictional research scenarios demonstrating four different possible designs for mixed method research.
@@ 5.5 Learning from Other Examples
So far, we have shared that using MMR enables SE researchers to study their research topics from a more holistic perspective, expanding and complimenting methods to achieve more and novel insights. We also note from the scenarios we provide (and the paper examples we reference in Section 4.2) that mixed methods are particularly valuable in studying socio-technical aspects of SE phenomena. The research scenarios we provided help to demonstrate the broad and varied landscape of possible MMR designs in SE research. The ACM SigSoft Empirical Standards for Multiple and Mixed Methods also provides a list of example MMR papers sigsoft , but care has to be taken in using these and other examples of MMR designs as “exemplars” to guide future work. With this in mind, we next present “antipatterns” to watch out for when designing or critiquing an MMR design. An earlier version of these antipatterns also appears in the same Standards, contributed by the first author of this paper.
6 Antipatterns of Mixed Method Research Designs
The principles and research design patterns we described above can be considered as useful guidelines to follow when designing a study. And through the scenarios above, we illustrated how the principles of MMR could be followed. In the following, we provide examples of “what not to do” and “what to avoid”.
As researchers, authors and reviewers of papers, we have noted a number of common antipatterns in how mixed methods are used or presented. Some of the antipatterns we encountered originate in issues with how the research is presented, rather than inherent to the conduct of the research. We introduce presentation antipatterns first noting how they can be potentially fixed. The other study design antipatterns are harder to “fix” after the research has been conducted, but hopefully can be spotted at design time (or during review of a paper). We recommend researchers think about these antipatterns early in their research process.